# Using Uncertainty Quantification to Characterize and Improve Out-of-Domain Learning for PDEs

### v1

@@ B.2 Diversity in the Heatmaps of the Ensemble
Appendix C Diversity Regularization Ablations
In our proposed DiverseNO, we solve the optimization problem in Equation 4. We test different regularizations in the second term of Equation 4 to enforce diversity in the last layer heads of DiverseNO: maximize the $L_{2}$ loss between (a) the weights of each head (ours), (b) the OOD predictions of each head, and (c) the gradients with respect to each head. For (b) and (c), we try two variants: the $L_{2}$ between the respective quantities after standardizing them to have mean 0 and variance 1, and without standardization. Note that (b) requires access to a set of OOD inputs without corresponding target outputs. Figures 17 -18 show that the proposed diversity measure results in best MSE and n-MeRCI performance across all OOD shifts. Regularizations that directly diversify the outputs or gradients are also sensitive to the regularization strength and require a careful trade-off between the prediction loss and regularization penalty. Proposed regularization is more robust to the hyperparameter and monotonically improves the performance for reasonable values of regularization strength.
@@ Appendix C Diversity Regularization Ablations
Figure 17: Diversity regularization ablation. Effect of different diversity regularizations on MSE for small, medium, large OOD shifts in the 1-d heat equation. Our proposed regularization of enforcing diversity over the weights of each head has the best performance across various regularization strengths and OOD shifts.
@@
Figure 18: Diversity regularization ablation. Effect of different diversity regularizations on n-MeRCI for small, medium, large OOD shifts in the 1-d heat equation. Our proposed regularization of enforcing diversity over the weights of each head has the best performance across various regularization strengths and OOD shifts.
@@ Appendix D PDE Test Problems
In this section, we provide the details of the various test problems that we study, and the construction of both the training and test OOD datasets (See Table 1 for a summary).
@@ D.1 Generalized Porous Medium Equation (GPME) Family
Linear advection Train $\beta\in[1,~{}~{}2]$
@@ D.1.1 Heat (Diffusion) Equation
The training dataset consists of $N=400$ input/output pairs $\{\phi^{(i)},u^{(i)}\}_{i=1}^{N}$ where $\phi^{(i)}(x,t):=k^{(i)},\forall x$, denotes a constant function over the domain representing the value of the diffusivity parameter and $u^{(i)}$ denotes the corresponding solutions at times $t\in[0,1]$. The $\phi^{(i)}$ are passed as input to the models without any normalization applied. During training, we consider the parameter $k^{(i)}\sim\text{Unif}(1,5)$, and evaluate the trained neural operator on small, medium and large OOD ranges for the diffusivity parameter: $k^{{\text{test}}}\sim\text{Unif}(5,6),\text{Unif}(6,7),\text{Unif}(7,8)$.
@@ D.1.2 Porous Medium Equation (PME)
We train the neural operator to map from a constant function denoting the degree $m$ that identifies $k(u)$ to the corresponding solution for all $t\in[0,1]$. The training dataset consists of $N=400$ input/output pairs $\{\phi^{(i)},u^{(i)}\}_{i=1}^{N}$ where $\phi^{(i)}(x,t):=m^{(i)},\forall x$, is a constant function over the domain representing the degree and $u^{(i)}$ denotes the corresponding solutions at times $t\in[0,1]$. The $\phi^{(i)}$ are passed as input to the models without any normalization applied. During training, we sample the parameter $m^{(i)}\sim\text{Unif}(2,3)$, and evaluate the trained neural operator on small, medium and large OOD ranges for the degree: $m^{{\text{test}}}\sim\text{Unif}(1,2),\text{Unif}(4,5),\text{Unif}(5,6)$. With increasing $m$, the solution becomes sharper and more challenging for the learned operator.
@@ D.1.3 Stefan Equation
We train the NO to map from a constant function denoting the parameter $u^{*}$ to the solution for all $t\in[0,0.1]$. The training dataset consists of $N=400$ input/output pairs $\{\phi^{(i)},u^{(i)}\}_{i=1}^{N}$ where $\phi^{(i)}(x,t)=u{{}^{*}}^{(i)},\forall x$, denotes a constant function over the domain representing the solution value at the shock and $u^{(i)}$ denotes the corresponding solutions at times $t\in[0,0.1]$. The $\phi^{(i)}$ are passed as input to the models without any normalization applied. During training, we sample the parameter $u{{}^{*}}^{(i)}\sim\text{Unif}(0.6,0.65)$, and evaluate the trained neural operator on small, medium and large OOD ranges: $u{{}^{*}}^{{\text{test}}}\sim\text{Unif}(0.55,0.6),\text{Unif}(0.7,0.75),\text{Unif}(0.5,0.55)$.
@@ D.2 Hyperbolic Linear Advection Equation
describes the motion of a fluid advected by a constant velocity $\beta>0$. We solve the PDE for initial conditions $u(x,0)={\bm{1}}_{x\leq 0.5},\forall x\in[0,1]$, and Dirichlet boundary conditions $u(0,t)=1,u(1,t)=0,\forall t$. The solution is a rightward moving shock (discontinuity) with the speed defined by the parameter $\beta$.
We train the neural operator to map from a constant function denoting the velocity parameter $\beta$ to the solution for all $t\in[0,1]$. During training, we sample the parameter $\beta^{(i)}\sim\text{Unif}(1,2)$, and evaluate the trained neural operator on small, medium and large OOD ranges for the $\beta$: $\beta^{{\text{test}}}\sim\text{Unif}(0.5,1),\text{Unif}(2.5,3),\text{Unif}(3,3.5)$.
@@ F.1 Detailed Metric Results and Solution Profiles
Figures 19-22 show the solution profiles for the “easy” smooth and parabolic heat equation with zero Dirichlet boundary conditions for in-domain, small, medium and large OOD shifts, respectively, of the diffusivity parameter $k$. We see that on this “easy” case, most methods are very accurate in-domain in Figure 19 and perform well in the small shift cases in Figure 20. Errors in the BayesianNO, VarianceNO and MC-DropoutNO baselines start to form for medium shifts in Figure 21 and grow in the large shift case in Figure 22. We see both EnsembleNO and DiverseNO output good uncertainty estimates (3 standard deviations) that contain the true solution within these error bounds. See corresponding metric results in Table 3.
Table 3: 1-d heat equation. MSE normal-$\downarrow$, NLL normal-$\downarrow$, n-MeRCI normal-$\downarrow$, RMSCE normal-$\downarrow$ and CRPS normal-$\downarrow$ (mean and standard deviation over 5 seeds) metrics for different UQ methods on the 1-d heat equation in-domain and with small, medium and large OOD shifts, where $k^{\text{train}}\in[1,5]$. Bold indicates values within one standard deviation of the best mean.
@@
Figures 23-26 show the solution profiles for the “medium” degenerate parabolic PME on an in-domain task and for small, medium and large OOD shifts, respectively, of the power $m$ in the monomial coefficient $k(u)=u^{m}$. The solution for larger values of $m\geq 1$ becomes sharper and more challenging. Figure 23 shows that all methods perform well on the in-domain task with the exception of VarianceNO and MC-DropoutNO having small negative oscillations at the sharp corner (degeneracy point), which separates the region with fluid to the left from the region without fluid to the right. Figure 24 shows the solutions on performing inference on an easier task where the values of $m$ are smaller than those trained on and the corresponding solution is smoother. We see in this easier case that all the methods perform reasonably well. There is left boundary error with the growing in time left boundary condition with VarianceNO, which grows as the OOD shift increases from medium in Figure 25 to large in Figure 26. The medium and large shift cases are particularly more challenging since we train on smaller values of $m$ and perform inference on the sharper cases with increased values $m$. As expected as the problems becomes harder and more challenging for larger $m$ (larger shifts), the uncertainty widens for EnsembleNO and DiverseNO. See corresponding metric results in Table 4.
Table 4: 1-d PME. MSE normal-$\downarrow$, NLL normal-$\downarrow$, n-MeRCI normal-$\downarrow$, RMSCE normal-$\downarrow$ and CRPS normal-$\downarrow$ (mean and standard deviation over 5 seeds) metrics for different UQ methods on the 1-d PME in-domain and with small, medium and large OOD shifts, where $m^{\text{train}}\in[2,3]$. Bold indicates values within one standard deviation of the best mean.
@@
Figures 27-30 show the solution profiles for the “hard” degenerate parabolic, discontinuous GPME, i.e., the Stefan equation, on an in-domain task and for small, medium and large OOD shifts, respectively, of the coefficient $u^{*}$ in $k(u)={\bm{1}}_{u\geq u^{*}}$ in Equation 9, where $u^{*}=u(x^{*}(t),t)$ denotes the solution value at the shock $x^{*}(t)$. The parameter $u^{*}$ also has an effect on the shock position, with smaller values of $u^{*}$ resulting in a faster shock speed. Figure 27 shows that the methods are relatively accurate on the in-domain task with EnsembleNO and DiverseNO exactly capturing the shock with tight uncertainty bounds. The small Gibbs phenomenon at the shock position that occurs with EnsembleNO is damped with DiverseNO. VarianceNO has large uncertainty around the diffused shock. We see that even for small OOD shifts in Figure 28 on this “hard” shock problem, several of the baseline methods suffer numerical artifacts of spurious oscillations, being overly diffusive and having an incorrect shock speed, where the predicted shock position either lags or is ahead of the true shock position, which worsens as the OOD shift increases in Figures 29-30. See corresponding metric results in Table 5.
Table 5: 1-d Stefan equation. MSE normal-$\downarrow$, NLL normal-$\downarrow$, n-MeRCI normal-$\downarrow$, RMSCE normal-$\downarrow$ and CRPS normal-$\downarrow$ (mean and standard deviation over 5 seeds) metrics for different UQ methods on the 1-d Stefan equation in-domain and with small, medium and large OOD shifts, where $u{{}^{*}}^{{\text{train}}}\in[0.6,0.65]$. Bold indicates values within one standard deviation of the best mean.

### v2

@@ B.2 Diversity in the Heatmaps of the Ensemble
Appendix C DiverseNO Ablations
In our proposed DiverseNO, we solve the following optimization problem in Equation 4 given as
$\hat{{\mathbf{\theta}}}=\operatorname*{arg\,min}_{{\mathbf{\theta}}}\underbrace{\frac{1}{NM}\sum_{i=1}^{N}\sum_{m=1}^{M}{\frac{||\hat{u}_{m}^{(i)}-u^{(i)}||^{2}_{L_{2}}}{||u^{(i)}||^{2}_{L_{2}}}}}_{\text{unconstrained NO loss}}\underbrace{-\frac{2\lambda_{\text{diverse}}}{M(M-1)}\sum_{m,k:m<k}||{\mathbf{\theta}}_{m}-{\mathbf{\theta}}_{k}||^{2}_{2}}_{\text{diversity regularization}}.$
In this section, we first study the effect of the hyperparameter $M$, which denotes the number of prediction heads on the accuracy. We then study the effect of various types of diversity regularizations.
C.1 Hyperparameter study on the number of prediction heads $M$
In this subsection, we perform a detailed hyperparameter study on the number of prediction heads $M$ across various metrics. We see that our choice of $M=10$ generally gives the lowest MSE (Figure 17), n-MeRCI (Figure 18) and number of FLOPs (Figure 19) for various FNO channel widths on in-domain and varying OOD tasks.
(a) In-domain
(b) OOD small
(c) OOD medium
(d) OOD large
Figure 17: Effect of $M$ on MSE. The MSE metric as a function of the number of prediction heads $M$ for various penultimate layer sizes. The diversity regularization strength $\lambda_{\text{diverse}}$ is fixed at 10. Increasing the number of prediction heads increases MSE mildly likely due to the fact that it is harder to train each of the prediction heads accurately to match the output. The MSE significantly decreases when increasing the FNO channel width, which is expected as the shared parameters of the model increase.
(a) OOD small
(b) OOD medium
(c) OOD large
Figure 18: Effect of $M$ on n-MeRCI. The n-MeRCI metric as a function of the the number of prediction heads $M$ for various penultimate layer sizes. The diversity regularization strength $\lambda_{\text{diverse}}$ is fixed at 10. The UQ estimates vary with the number of prediction heads. A large value of $M$ generally degrades the n-MeRCI metric given a fixed regularization strength because we cannot diversify all possible pairs of prediction heads as it is computationally expensive.
Figure 19: Effect of $M$ on FLOPs. The number of FLOPs for one forward pass as a function of the number of prediction heads $M$ for various penultimate layer sizes. The diversity regularization strength $\lambda_{\text{diverse}}$ is fixed at 10. The number of FLOPs increases mildly with the number of prediction heads.
C.2 Diversity Regularization Ablations
We test different regularizations in the second term of Appendix C to enforce diversity in the last layer heads of DiverseNO: maximize the $L_{2}$ loss between (a) the weights of each head (ours), (b) the OOD predictions of each head, and (c) the gradients with respect to each head. For (b) and (c), we try two variants: the $L_{2}$ between the respective quantities after standardizing them to have mean 0 and variance 1, and without standardization. Note that (b) requires access to a set of OOD inputs without corresponding target outputs. Figures 20-21 show that the proposed diversity measure results in best MSE and n-MeRCI performance across all OOD shifts. Regularizations that directly diversify the outputs or gradients are also sensitive to the regularization strength and require a careful trade-off between the prediction loss and regularization penalty. Our proposed regularization is more robust to the diversity hyperparameter $\lambda_{\text{diverse}}$, and monotonically improves the performance for reasonable values of this regularization strength.
The MSE (Figure 22) and especially the n-MeRCI metric (Figure 23) monotonically improve (decrease) as a function of $\lambda_{\text{diverse}}$ with larger values of this regularization strength $\approx 10$ being favorable. These figures also show that without diversity regularization, i.e., $\lambda_{\text{diverse}}=0$, there is a large gap in performance between DiverseNO and EnsembleNO, which is closed with larger $\lambda_{\text{diverse}}$.
@@ C.2 Diversity Regularization Ablations
Figure 20: MSE diversity regularization ablation. Effect of different diversity regularizations on MSE for small, medium, large OOD shifts in the 1-d heat equation. Our proposed regularization of enforcing diversity over the weights of each head has the best performance across various regularization strengths and OOD shifts.
@@
Figure 21: n-MeRCI diversity regularization ablation. Effect of different diversity regularizations on n-MeRCI for small, medium, large OOD shifts in the 1-d heat equation. Our proposed regularization of enforcing diversity over the weights of each head has the best performance across various regularization strengths and OOD shifts.
(a) Small OOD shift
(b) Medium OOD shift
(c) Large OOD shift
Figure 22: Effect of $\lambda_{\text{diverse}}$ on MSE. Comparison of the MSE metric as a function of the diversity regularization strength $\lambda_{\text{diverse}}$ for DiverseNO to the MSE of the other baselines.
(a) Small OOD shift
(b) Medium OOD shift
(c) Large OOD shift
Figure 23: Effect of $\lambda_{\text{diverse}}$ on n-MeRCI. Comparison of the n-MeRCI metric as a function of the diversity regularization strength $\lambda_{\text{diverse}}$ for DiverseNO to the n-MeRCI of the other baselines.
@@ Appendix D PDE Test Problems
In this section, we provide the details of the various test problems that we study, and the construction of both the training and test OOD datasets. See Table 1 for a summary.
@@ D.1 Generalized Porous Medium Equation (GPME) Family
Linear advection, constant input Train $\beta\in[1,~{}~{}2]$
@@
Linear advection, non-constant input Train $a\in[0.45,0.55]$
$u_{t}+u_{x}=0$, OOD small $a\in[0.4,0.45]$
$u(x,0)={\bm{1}}_{x\leq a}$ OOD medium $a\in[0.6,0.65]$
OOD large $a\in[0.35,0.4]$
2-d Darcy Flow Train $k\in[3,4]$
$-\nabla\cdot(k\nabla u(x))=1$ OOD small $k\in[4,4.5]$
OOD medium $k\in[4.5,5]$
OOD large $k\in[5,6]$
@@ D.1.1 Heat (Diffusion) Equation
The training dataset consists of $N=400$ input/output pairs $\{\phi^{(i)},u^{(i)}\}_{i=1}^{N}$ where $\phi^{(i)}(x,t):=k^{(i)},\forall x$, denotes a constant function over the domain representing the value of the diffusivity parameter and $u^{(i)}$ denotes the corresponding solutions at times $t\in[0,1]$. The $\phi^{(i)}$ are passed as input to the models without any normalization applied. During training, we consider the parameter $k^{(i)}\sim\text{Unif}(1,5)$, and evaluate the trained NO on small, medium and large OOD ranges for the diffusivity parameter: $k^{{\text{test}}}\sim\text{Unif}(5,6),\text{Unif}(6,7),\text{Unif}(7,8)$, respectively.
@@ D.1.2 Porous Medium Equation (PME)
We train the NO to map from a constant function denoting the degree $m$ that identifies $k(u)$ to the corresponding solution for all $t\in[0,1]$. The training dataset consists of $N=400$ input/output pairs $\{\phi^{(i)},u^{(i)}\}_{i=1}^{N}$ where $\phi^{(i)}(x,t):=m^{(i)},\forall x$, is a constant function over the domain representing the degree and $u^{(i)}$ denotes the corresponding solutions at times $t\in[0,1]$. The $\phi^{(i)}$ are passed as input to the models without any normalization applied. During training, we sample the parameter $m^{(i)}\sim\text{Unif}(2,3)$, and evaluate the trained NO on small, medium and large OOD ranges for the degree: $m^{{\text{test}}}\sim\text{Unif}(1,2),\text{Unif}(4,5),\text{Unif}(5,6)$, respectively. With increasing $m$, the solution becomes sharper and more challenging for the learned operator.
@@ D.1.3 Stefan Equation
We train the NO to map from a constant function denoting the parameter $u^{*}$ to the solution for all $t\in[0,0.1]$. The training dataset consists of $N=400$ input/output pairs $\{\phi^{(i)},u^{(i)}\}_{i=1}^{N}$ where $\phi^{(i)}(x,t)=u{{}^{*}}^{(i)},\forall x$, denotes a constant function over the domain representing the solution value at the shock and $u^{(i)}$ denotes the corresponding solutions at times $t\in[0,0.1]$. The $\phi^{(i)}$ are passed as input to the models without any normalization applied. During training, we sample the parameter $u{{}^{*}}^{(i)}\sim\text{Unif}(0.6,0.65)$, and evaluate the trained NO on small, medium and large OOD ranges: $u{{}^{*}}^{{\text{test}}}\sim\text{Unif}(0.55,0.6),\text{Unif}(0.7,0.75),\text{Unif}(0.5,0.55)$, respectively.
@@ D.2 Hyperbolic Linear Advection Equation
describes the motion of a fluid advected by a constant velocity $\beta>0$. We consider the following two tasks for the PDE.
Constant parameter to solution mapping.
We solve the PDE for initial conditions $u(x,0)={\bm{1}}_{x\leq 0.5},\forall x\in[0,1]$, and Dirichlet boundary conditions $u(0,t)=1,u(1,t)=0,\forall t$. The solution is a rightward moving shock (discontinuity) with the speed defined by the parameter $\beta$ (Hansen et al., 2023).
We train the NO to map from a constant function denoting the velocity parameter $\beta$ to the solution for all $t\in[0,0.1]$. The training dataset consists of $N=400$ input/output pairs $\{\phi^{(i)},u^{(i)}\}_{i=1}^{N}$ where $\phi^{(i)}(x,t)=\beta^{(i)},\forall x$, denotes a constant function over the domain representing the velocity and $u^{(i)}$ denotes the corresponding solutions at times $t\in[0,0.1]$. During training, we sample the parameter $\beta^{(i)}\sim\text{Unif}(1,2)$, and evaluate the trained NO on small, medium and large OOD ranges for the $\beta$: $\beta^{{\text{test}}}\sim\text{Unif}(0.5,1),\text{Unif}(2.5,3),\text{Unif}(3,3.5)$, respectively.
Non-constant initial condition to solution mapping.
We solve the PDE for various initial conditions $u(x,0)={\bm{1}}_{x\leq a},\forall x\in[0,1]$, where $a$ denotes the initial shock location. We use Dirichlet boundary conditions $u(0,t)=1,u(1,t)=0,\forall t$, and a fixed speed $\beta=1$.
We train the NO to map from the non-constant function denoting the initial condition, i.e., $u(x,0)$ to the solution for all $t\in[0,0.1]$. The training dataset consists of $N=400$ input/output pairs $\{\phi^{(i)},u^{(i)}\}_{i=1}^{N}$ where $\phi^{(i)}(x,t)=u^{(i)}(x,0),\forall x$, denotes a constant function over the domain representing the velocity and $u^{(i)}$ denotes the corresponding solutions at times $t\in[0,0.1]$. During training, we generate initial conditions using shock locations $a^{(i)}\sim\text{Unif}(0.45,0.55)$, and evaluate the trained NO on small, medium and large OOD ranges for the $a$: $a^{{\text{test}}}\sim\text{Unif}(0.4,0.45),\text{Unif}(0.6,0.65),\text{Unif}(0.35,0.4)$, respectively.
D.3 2-d Elliptic Darcy Flow
We consider the steady state solution of the 2-d Darcy flow equation (linear, elliptic):
$-\nabla\cdot(k\nabla u(x))=f(x)\>,\qquad x\in[0,1]^{2},$
with Dirichlet boundary conditions $u(x)=0$ for all $x$ on the boundary, forcing function $f(x)=1$, and permeability field defined by the parameter $k$.
We train the NO to map from the constant scalar field over the 2-d domain denoting the permeability $k$ of the surface to the steady-state solution. The training dataset consists of $N=400$ input/output pairs $\{\phi^{(i)},u^{(i)}\}_{i=1}^{N}$ where $\phi^{(i)}(x)=k^{(i)},\forall x$, denotes a constant field over the domain representing the permeability and and $u^{(i)}$ denotes the corresponding solutions representing the unknown pressure. During training, we sample the permeability parameter $k^{(i)}\sim\text{Unif}(3,4)$, and evaluate the trained NO on small, medium and large OOD ranges for the $k$: $k^{{\text{test}}}\sim\text{Unif}(4,4.5),\text{Unif}(4.5,5),\text{Unif}(5,6)$, respectively.
@@ Appendix E Detailed Experiment Settings
We use the standard optimization procedure for training FNO models (Li et al., 2020a). In particular, we use the Adam optimizer with a weight decay. We optimize the objective, and learn over batches of a given batch size $B$ (fixed to $B=20$ in our experiments). We use a learning rate scheduler that halves the learning rate after every 50 epochs.
@@ F.1 Detailed Metric Results and Solution Profiles
Figures 24-27 show the solution profiles for the “easy” smooth and parabolic heat equation with zero Dirichlet boundary conditions for in-domain, small, medium and large OOD shifts, respectively, of the diffusivity parameter $k$. We see that on this “easy” case, most methods are very accurate in-domain in Figure 24 and perform well in the small shift cases in Figure 25. Errors in the BayesianNO, VarianceNO and MC-DropoutNO baselines start to form for medium shifts in Figure 26 and grow in the large shift case in Figure 27. We see both EnsembleNO and DiverseNO output good uncertainty estimates (3 standard deviations) that contain the true solution within these error bounds. See corresponding metric results in Table 3.
Table 3: 1-d heat equation. MSE $\downarrow$, NLL $\downarrow$, n-MeRCI $\downarrow$, RMSCE $\downarrow$ and CRPS $\downarrow$ (mean and standard deviation over 5 seeds) metrics for different UQ methods on the 1-d heat equation in-domain and with small, medium and large OOD shifts, where $k^{\text{train}}\in[1,5]$. Bold indicates values within one standard deviation of the best mean.
@@
Figures 28-31 show the solution profiles for the “medium” degenerate parabolic PME on an in-domain task and for small, medium and large OOD shifts, respectively, of the power $m$ in the monomial coefficient $k(u)=u^{m}$. The solution for larger values of $m\geq 1$ becomes sharper and more challenging. Figure 28 shows that all methods perform well on the in-domain task with the exception of VarianceNO and MC-DropoutNO having small negative oscillations at the sharp corner (degeneracy point), which separates the region with fluid to the left from the region without fluid to the right. Figure 29 shows the solutions on performing inference on an easier task where the values of $m$ are smaller than those trained on and the corresponding solution is smoother. We see in this easier case that all the methods perform reasonably well. There is left boundary error with the growing in time left boundary condition with VarianceNO, which grows as the OOD shift increases from medium in Figure 30 to large in Figure 31. The medium and large shift cases are particularly more challenging since we train on smaller values of $m$ and perform inference on the sharper cases with increased values $m$. As expected as the problems becomes harder and more challenging for larger $m$ (larger shifts), the uncertainty widens for EnsembleNO and DiverseNO. See corresponding metric results in Table 4.
Table 4: 1-d PME. MSE $\downarrow$, NLL $\downarrow$, n-MeRCI $\downarrow$, RMSCE $\downarrow$ and CRPS $\downarrow$ (mean and standard deviation over 5 seeds) metrics for different UQ methods on the 1-d PME in-domain and with small, medium and large OOD shifts, where $m^{\text{train}}\in[2,3]$. Bold indicates values within one standard deviation of the best mean.
@@
Figures 32-35 show the solution profiles for the “hard” degenerate parabolic, discontinuous GPME, i.e., the Stefan equation, on an in-domain task and for small, medium and large OOD shifts, respectively, of the coefficient $u^{*}$ in $k(u)={\bm{1}}_{u\geq u^{*}}$ in Equation 9, where $u^{*}=u(x^{*}(t),t)$ denotes the solution value at the shock $x^{*}(t)$. The parameter $u^{*}$ also has an effect on the shock position, with smaller values of $u^{*}$ resulting in a faster shock speed. Figure 32 shows that the methods are relatively accurate on the in-domain task with EnsembleNO and DiverseNO exactly capturing the shock with tight uncertainty bounds. The small Gibbs phenomenon at the shock position that occurs with EnsembleNO is damped with DiverseNO. VarianceNO has large uncertainty around the diffused shock. We see that even for small OOD shifts in Figure 33 on this “hard” shock problem, several of the baseline methods suffer numerical artifacts of spurious oscillations, being overly diffusive and having an incorrect shock speed, where the predicted shock position either lags or is ahead of the true shock position, which worsens as the OOD shift increases in Figures 34-35. See corresponding metric results in Table 5.
Table 5: 1-d Stefan equation. MSE $\downarrow$, NLL $\downarrow$, n-MeRCI $\downarrow$, RMSCE $\downarrow$ and CRPS $\downarrow$ (mean and standard deviation over 5 seeds) metrics for different UQ methods on the 1-d Stefan equation in-domain and with small, medium and large OOD shifts, where $u{{}^{*}}^{{\text{train}}}\in[0.6,0.65]$. Bold indicates values within one standard deviation of the best mean.